OpenStudy (asad786):

Which is an example of the opposite of a sum property? A. –17 + 32 = 32 + (–17) B. –26 + (4 + 8) = (–26 + 4) + 8 C. –5(6 + 3) = –5(6) + –5(3) D. –(–17 + 5) = 17 + (–5)

OpenStudy (asad786):

Which is an example of the opposite of a sum property? A. –17 + 32 = 32 + (–17) B. –26 + (4 + 8) = (–26 + 4) + 8 C. –5(6 + 3) = –5(6) + –5(3) D. –(–17 + 5) = 17 + (–5)

OpenStudy (anonymous):

it is D @Asad786